Cans of regular soda have volumes with a mean of 13.77 oz and a standard deviation of 0.12 oz. Is it unusual for a can to contain 14.17 oz of soda?
3. (15) Cans of regular soda have volumes with a mean of 13.77 oz and a standard deviation Of 0.12 oz. Is it unusual for a can to contain 14.17 oz of soda? A. Minimum usual value B. Maximum usual value C. Is 14.17 oz an unusual volume? (Circle the correct OZ oZ response.) No, because it is larger than the maximum usual value. Yes, because it is larger than the maximum usual value Yes, because it is between the minimum and maximum usual value:s Yes, because it is smaller than the minimum usual value No, because it is between the minimum and maximum usual values No, because it is smaller than the minimum usual value
Answer #1

According to empirical property of distributions,

99% values lie between (mean - 3*sd, mean + 3*sd)

Here,

mean - 3*standard deviation = 13.77 - (3*0.12) = 13.41

mean + 3*standard deviation = 13.77 + (3*0.12) = 14.13

Since 14.17 lies outside the interval (13.41, 14.13), so we can say it is unusual for a can to contain 14.17 oz of soda.

a) Minimum usual value = 13.41

b) Maximum usual value = 14.13

c) Yes, because it is larger than the maximum "usual" value.